### v1

@@ 3. $(\mathcal{S}_{3})$
      Then $F^{c}=\{m+1,m+4\ (\text{mod $n$}),n-1\}$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<n-1$, we get $n\geq 12$. Further, if $n\in\{12,13\}$, then $n-1=m+5$, and thus, $F^{c}=\{m+1,(m+1)+3,(m+1)+4\}$. It follows from Proposition 3.14 that $F$ is not a spanning facet. This is a contradiction. So, assume that $n\geq 14$. Then $p_{1}=m+4\ (\text{mod $n$})=m+4$ and $p_{1}+1=m+5<n-1$. Thus $p_{1}+1\in F$. Let $F^{\prime c}=\{m+1\}\sqcup\{m+4,m+5\}$. Then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and satisfies $(\mathcal{D}_{4})$. Hence $F^{\prime}\in\mathcal{D}$. Using Proposition 3.18, it follows that $F$ is not a spanning facet, a contradiction. So $p_{1}\neq m+4\ (\text{mod $n$})$
@@
    We show that $F$ satisfies $(\mathcal{S}_{3})$. Since $\min\{2m-\alpha_{p},\alpha_{p}-4\}=2m-\alpha_{p}$, we prove that $p_{1}\in V(W_{n})\setminus(\{2m-\alpha_{p},\ldots,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\})$, where
@@
    We have $\alpha_{p}\geq m+2$ and $\alpha_{p}<_{\mathcal{O}}p_{1}$. Thus, either $p_{1}<2m-\alpha_{p}$ or $p_{1}>\alpha_{p}$ by Remark 3.1 (ii). This means that $p_{1}\notin\{2m-\alpha_{p},\ldots,\alpha_{p}-1,\alpha_{p}\}$. Therefore, it is sufficient to show that $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\}$.
@@
      Here, since $\alpha_{p}\leq n-6<n-4$, we have $y=n-1$. Clearly, $p_{1}\neq y$. Suppose $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+2\}$. Then $p_{1}+1\leq\alpha_{p}+3\leq n-3<n-1$, and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Further, since $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=W_{n}[\{\alpha_{p},p_{1},p_{1}+1\}]$ is connected, $F$ is not a spanning facet by Proposition 3.18. This is a contradiction. Therefore $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2\}$.
      Suppose $p_{1}=\alpha_{p}+3\text{ (mod \ $n$})$. We have $\alpha_{p}+3\leq n-3<n-1$. So $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Then $p_{1}+1=\alpha_{p}+4\leq n-2<n-1$, and thus, $p_{1}+1\in F$. Clearly, $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. If $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and satisfies $(\mathcal{D}_{4})$. Thus $F^{\prime}\in\mathcal{D}$. By Proposition 3.18, it follows that $F$ is not a spanning facet, a contradiction. Therefore, $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
@@
      We have $y=n-1$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Further, $\alpha_{p}+4=n-1=p_{2}$. Therefore, since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+2,y\}$. Moreover, if $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+3\}$, then either $F^{c}=\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+4\}$ or $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$. Then, by Proposition 3.14, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
      Then $y=0$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3=n-1=p_{2}$. If $p_{1}=y$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\ (\text{mod \ $n$})\}$, which contradicts the fact that $F$ is a spanning facet by Proposition 3.14. Hence $p_{1}\neq y$. Further, since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, we get $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
      We have $y=1$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$}),1)\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
In Section 3.1, we have proved that $\prec$ gives a shelling order for $\Delta_{3}(W_{n})$. Hence the $3$-cut complex $\Delta_{3}(W_{n})$ is shellable for $n\geq 9$.
Now, using Lemmas 3.13 and 3.19, we see that $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet for the shelling order given by $\prec$ on $\Delta_{3}(W_{n})$ if and only if $F\in\mathcal{S}$. Therefore, to count the number of spanning facets, it is enough to count the total facets in $\mathcal{S}$. Let $F\in\mathcal{S}$. Then $F\in\mathcal{T}_{s}$ for some $s$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and $F$ satisfies one of the conditions from $(\mathcal{S}_{1})$ to $(\mathcal{S}_{3})$.
@@
Suppose $F$ satisfies $(\mathcal{S}_{3})$. If $\alpha_{s}\in\{m-1,m,m+1\}$, then $\omega=\alpha_{s}-4$. For each $\alpha_{s}$, there are $n-9$ possibilities for $s_{1}$. Therefore, we have $3(n-9)$ spanning facets. If $\alpha_{s}\in\{m+2,\ldots,n-3\}$, then $\omega=2m-\alpha_{s}$. There are $n-m-4$ possibilities for $\alpha_{s}$, and for each value of $\alpha_{s}$, there are $n-(2\alpha_{s}-2m+5)$ possibilities for $s_{1}$. Thus, the number of spanning facets satisfying $(\mathcal{S}_{3})$
$=(n-m-4)(n+2m-5)-\left(2\sum_{\alpha_{s}=m+2}^{n-3}\alpha_{s}\right)$
$=(n-m-4)(n+2m-5)-(n-3)(n-2)+(m+1)(m+2)$
$=mn-m^{2}-4n+16.$
@@
Hence from Theorem 2.6, we get $\Delta_{3}(W_{n})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$

### v2

@@ 3. $(\mathcal{S}_{3})$
      Then $F^{c}=\{m+1,m+4\ (\text{mod $n$}),n-1\}$. Since $C_{n}^{2}[F^{c}]$ is disconnected and $p_{1}<n-1$, we get $n\geq 12$. Further, if $n\in\{12,13\}$, then $n-1=m+5$, and thus, $F^{c}=\{m+1,(m+1)+3,(m+1)+4\}$. It follows from Proposition 3.15 that $F$ is not a spanning facet. This is a contradiction. So, assume that $n\geq 14$. Then $p_{1}=m+4\ (\text{mod $n$})=m+4$ and $p_{1}+1=m+5<n-1$. Thus $p_{1}+1\in F$. Let $F^{\prime c}:=\{m+1\}\sqcup\{m+4,m+5\}$. Then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ and satisfies $(\mathcal{D}_{4})$. Hence $F^{\prime}\in\mathcal{D}$. Using Proposition 3.19, it follows that $F$ is not a spanning facet, a contradiction. So $p_{1}\neq m+4\ (\text{mod $n$})$
@@
    We show that $F$ satisfies $(\mathcal{S}_{3})$. Since $\min\{2m-\alpha_{p},\alpha_{p}-4\}=2m-\alpha_{p}$, we prove that $p_{1}\in V(C_{n}^{2})\setminus(\{2m-\alpha_{p},\ldots,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\})$, where
@@
    We have $\alpha_{p}\geq m+2$ and $\alpha_{p}<_{\mathcal{O}}p_{1}$. Thus, either $p_{1}<2m-\alpha_{p}$ or $p_{1}>\alpha_{p}$ by Remark 3.1 (ii). This means that $p_{1}\notin\{2m-\alpha_{p},\ldots,\alpha_{p}-1,\alpha_{p}\}$. Hence, it is sufficient to show that $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\}$.
@@
      Here, since $\alpha_{p}\leq n-6<n-4$, we have $y=n-1$. Clearly, $p_{1}\neq y$. Suppose $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+2\}$. Then $p_{1}+1\leq\alpha_{p}+3\leq n-3<n-1$, and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Further, since $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=C_{n}^{2}[\{\alpha_{p},p_{1},p_{1}+1\}]$ is connected, $F$ is not a spanning facet by Proposition 3.19. This is a contradiction. Therefore $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2\}$.
      Suppose $p_{1}=\alpha_{p}+3\text{ (mod \ $n$})$. We have $\alpha_{p}+3\leq n-3<n-1$. So $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Then $p_{1}+1=\alpha_{p}+4\leq n-2<n-1$, and thus, $p_{1}+1\in F$. Clearly, $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. If $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ and satisfies $(\mathcal{D}_{4})$. Thus $F^{\prime}\in\mathcal{D}$. By Proposition 3.19, it follows that $F$ is not a spanning facet, a contradiction. Therefore, $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
@@
      We have $y=n-1$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Further, $\alpha_{p}+4=n-1=p_{2}$. Therefore, since $C_{n}^{2}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+2,y\}$. Moreover, if $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+3\}$, then either $F^{c}=\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+4\}$ or $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$. Then, by Proposition 3.15, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
      Then $y=0$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3=n-1=p_{2}$. If $p_{1}=y$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\ (\text{mod \ $n$})\}$, which contradicts the fact that $F$ is a spanning facet by Proposition 3.15. Hence $p_{1}\neq y$. Further, since $C_{n}^{2}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, we get $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
      We have $y=1$. Since $C_{n}^{2}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$}),1)\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
In Section 3.1, we have proved that $\prec$ gives a shelling order for $\Delta_{3}(C_{n}^{2})$. Hence the $3$-cut complex $\Delta_{3}(C_{n}^{2})$ is shellable for $n\geq 9$.
Now, using Lemmas 3.14 and 3.20, we see that $F\in M(\Delta_{3}(C_{n}^{2}))$ is a spanning facet for the shelling order given by $\prec$ on $\Delta_{3}(C_{n}^{2})$ if and only if $F\in\mathcal{S}$. Therefore, to count the number of spanning facets, it is enough to count the total facets in $\mathcal{S}$. Let $F\in\mathcal{S}$. Then $F\in\mathcal{T}_{s}$ for some $s$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and $F$ satisfies one of the conditions from $(\mathcal{S}_{1})$ to $(\mathcal{S}_{3})$.
@@
Suppose $F$ satisfies $(\mathcal{S}_{3})$. If $\alpha_{s}\in\{m-1,m,m+1\}$, then $\omega=\alpha_{s}-4$. For each $\alpha_{s}$, there are $n-9$ possibilities for $s_{1}$. If $\alpha_{s}\in\{m+2,\ldots,n-3\}$, then $\omega=2m-\alpha_{s}$. There are $n-m-4$ possibilities for $\alpha_{s}$, and for each value of $\alpha_{s}$, there are $n-(2\alpha_{s}-2m+5)$ possibilities for $s_{1}$. Thus, the number of spanning facets satisfying $(\mathcal{S}_{3})$
$=3(n-9)+(n-m-4)(n+2m-5)-\left(2\sum_{\alpha_{s}=m+2}^{n-3}\alpha_{s}\right)$
$=3(n-9)+(n-m-4)(n+2m-5)-(n-3)(n-2)+(m+1)(m+2)$
$=mn-m^{2}-n-11.$
@@
Hence from Theorem 2.5, we get $\Delta_{3}(C_{n}^{2})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$
4. Conclusion and Future Directions
In [1] and [2], the authors completely determined the homotopy type of the complex $\Delta_{k}(C_{n})$ for all $k$. They also proved that the complex $\Delta_{2}(C_{n}^{2})$ (if nonvoid) is not shellable. We have proved that the complex $\Delta_{3}(C_{n}^{2})$ of the squared cycle graphs is shellable for all $n\geq 9$, by constructing a shelling order. We believe that this shelling order can be generalized to prove that the complex $\Delta_{k}(C_{n}^{2})$ is shellable for all $k\geq 4$ and $n\geq k+6$ (1.3).
We further investigated the $k$-cut complex $\Delta_{k}(C_{n}^{p})$ for powered cycle graphs $C_{n}^{p}$ and, using SageMath, computed their homology groups (with coefficient $\mathbb{Z}$). We have data for some small values of $k$, $p$ and $n$, which is illustrated in Tables 1-5. Missing entries in these tables indicate that the corresponding complex $\Delta_{k}(C_{n}^{p})$ is void. An entry of the form $i:\mathbb{Z}^{\beta}$ denotes that the $i$-th homology group is $\mathbb{Z}^{\beta}$.
$p$ $n$ 8 9 10 11 12 13 14 15 16
3 $2:\mathbb{Z}$ $4:\mathbb{Z}^{2}$ $6:\mathbb{Z}$ $7:\mathbb{Z}$ $8:\mathbb{Z}$ $9:\mathbb{Z}$ $10:\mathbb{Z}$ $11:\mathbb{Z}$ $12:\mathbb{Z}$
4 $3:\mathbb{Z}$ $5:\mathbb{Z}$ $7:\mathbb{Z}^{3}$ $9:\mathbb{Z}$ $10:\mathbb{Z}$ $11:\mathbb{Z}$ $12:\mathbb{Z}$
5 $4:\mathbb{Z}$ $7:\mathbb{Z}$ $8:\mathbb{Z}$ $10:\mathbb{Z}^{4}$ $12:\mathbb{Z}$
6 $5:\mathbb{Z}$ $8:\mathbb{Z}^{2}$ $10:\mathbb{Z}$
Table 1. Homology of $2$-cut complexes $\Delta_{2}(C_{n}^{p})$
Note that the nonvoid $2$-cut complexes $\Delta_{2}(C_{n}^{p})$ in Table 1, have non trivial homology in dimension lower than the dimension of $\Delta_{2}(C_{n}^{p})$. We propose the following conjecture.
Conjecture 4.1.
For $p\geq 3$, if $\Delta_{2}(C_{n}^{p})\neq\emptyset$, then $\Delta_{2}(C_{n}^{p})$ is not shellable.
A natural question arises here: What is the homotopy type of $\Delta_{2}(C_{n}^{p})$ for $p\geq 3$?
$n$ $p$ 2 3 4 5 6
7 $1:\mathbb{Z}$
8 $3:\mathbb{Z}$
9 $5:\mathbb{Z}$ $3:\mathbb{Z}$
10 $6:\mathbb{Z}^{6}$ $4:\mathbb{Z}^{11}$
11 $7:\mathbb{Z}^{12}$ $5:\mathbb{Z}$ $3:\mathbb{Z}$
12 $8:\mathbb{Z}^{19}$ $7:\mathbb{Z}^{2}$ $6:\mathbb{Z}^{6}$
13 $9:\mathbb{Z}^{27}$ $9:\mathbb{Z}$ $7:\mathbb{Z}^{40}$ $5:\mathbb{Z}$
14 $10:\mathbb{Z}^{36}$ $10:\mathbb{Z}^{8}$ $8:\mathbb{Z}^{15}$ $6:\mathbb{Z}$
15 $11:\mathbb{Z}^{46}$ $11:\mathbb{Z}^{16}$ $9:\mathbb{Z}$ $9:\mathbb{Z}^{21}$ $5:\mathbb{Z}$
16 $12:\mathbb{Z}^{57}$ $12:\mathbb{Z}^{25}$ $11:\mathbb{Z}^{3}$ $10:\mathbb{Z}^{101}$ $8:\mathbb{Z}^{31}$
17 $13:\mathbb{Z}^{69}$ $13:\mathbb{Z}^{35}$ $13:\mathbb{Z}$ $11:\mathbb{Z}^{52}$ $9:\mathbb{Z}$
18 $14:\mathbb{Z}^{82}$ $14:\mathbb{Z}^{46}$ $14:\mathbb{Z}^{10}$ $12:\mathbb{Z}^{19}$ $12:\mathbb{Z}^{55}$
19 $15:\mathbb{Z}^{96}$ $15:\mathbb{Z}^{58}$ $15:\mathbb{Z}^{20}$ $13:\mathbb{Z}$ $13:\mathbb{Z}^{210}$
20 $16:\mathbb{Z}^{111}$ $16:\mathbb{Z}^{71}$ $16:\mathbb{Z}^{31}$ $15:\mathbb{Z}^{4}$ $14:\mathbb{Z}^{126}$
21 $17:\mathbb{Z}^{127}$ $17:\mathbb{Z}^{85}$ $17:\mathbb{Z}^{43}$ $17:\mathbb{Z}$ $15:\mathbb{Z}^{64}$
22 $18:\mathbb{Z}^{144}$ $18:\mathbb{Z}^{100}$ $18:\mathbb{Z}^{56}$ $18:\mathbb{Z}^{12}$ $16:\mathbb{Z}^{23}$
23 $19:\mathbb{Z}^{162}$ $19:\mathbb{Z}^{116}$ $19:\mathbb{Z}^{70}$ $19:\mathbb{Z}^{24}$ $17:\mathbb{Z}$
24 $20:\mathbb{Z}^{181}$ $20:\mathbb{Z}^{133}$ $20:\mathbb{Z}^{85}$ $20:\mathbb{Z}^{37}$ $19:\mathbb{Z}^{5}$
25 $21:\mathbb{Z}^{201}$ $21:\mathbb{Z}^{151}$ $21:\mathbb{Z}^{101}$ $21:\mathbb{Z}^{51}$ $21:\mathbb{Z}$
Table 2. Homology of $3$-cut complexes $\Delta_{3}(C_{n}^{p})$
$n$ $p$ 2 3 4 5 6
8 $1:\mathbb{Z}$
9 $3:\mathbb{Z}$
$4:\mathbb{Z}^{3}$
10 $5:\mathbb{Z}^{16}$ $2:\mathbb{Z}$
$3:\mathbb{Z}^{2}$
11 $6:\mathbb{Z}^{43}$ $4:\mathbb{Z}^{21}$
12 $7:\mathbb{Z}^{81}$ $5:\mathbb{Z}^{4}$ $3:\mathbb{Z}$
$6:\mathbb{Z}^{2}$ $4:\mathbb{Z}^{3}$
$7:\mathbb{Z}^{4}$
13 $8:\mathbb{Z}^{129}$ $7:\mathbb{Z}^{14}$ $5:\mathbb{Z}^{27}$
$8:\mathbb{Z}^{13}$ $6:\mathbb{Z}^{13}$
14 $9:\mathbb{Z}^{188}$ $8:\mathbb{Z}^{15}$ $7:\mathbb{Z}^{99}$ $4:\mathbb{Z}$
$9:\mathbb{Z}^{35}$ $5:\mathbb{Z}^{2}$
15 $10:\mathbb{Z}^{259}$ $9:\mathbb{Z}^{15}$ $8:\mathbb{Z}^{77}$ $6:\mathbb{Z}^{5}$
$10:\mathbb{Z}^{82}$ $10:\mathbb{Z}^{5}$ $7:\mathbb{Z}^{31}$
$8:\mathbb{Z}^{3}$
16 $11:\mathbb{Z}^{343}$ $10:\mathbb{Z}^{16}$ $9:\mathbb{Z}^{11}$ $8:\mathbb{Z}^{181}$ $5:\mathbb{Z}^{2}$
$11:\mathbb{Z}^{151}$ $10:\mathbb{Z}^{8}$ $9:\mathbb{Z}^{48}$ $6:\mathbb{Z}$
$11:\mathbb{Z}^{16}$
17 $12:\mathbb{Z}^{441}$ $11:\mathbb{Z}^{17}$ $11:\mathbb{Z}^{52}$ $9:\mathbb{Z}^{18}$ $7:\mathbb{Z}$
$12:\mathbb{Z}^{237}$ $12:\mathbb{Z}^{34}$ $10:\mathbb{Z}^{272}$ $8:\mathbb{Z}^{51}$
18 $13:\mathbb{Z}^{554}$ $12:\mathbb{Z}^{18}$ $12:\mathbb{Z}^{73}$ $11:\mathbb{Z}^{323}$ $9:\mathbb{Z}^{16}$
$13:\mathbb{Z}^{338}$ $13:\mathbb{Z}^{69}$ $13:\mathbb{Z}^{6}$ $10:\mathbb{Z}^{159}$
$11:\mathbb{Z}^{18}$
19 $14:\mathbb{Z}^{683}$ $13:\mathbb{Z}^{19}$ $13:\mathbb{Z}^{77}$ $12:\mathbb{Z}^{189}$ $11:\mathbb{Z}^{723}$
$14:\mathbb{Z}^{455}$ $14:\mathbb{Z}^{133}$ $14:\mathbb{Z}^{19}$ $12:\mathbb{Z}^{133}$
20 $15:\mathbb{Z}^{829}$ $14:\mathbb{Z}^{20}$ $14:\mathbb{Z}^{80}$ $13:\mathbb{Z}^{28}$ $12:\mathbb{Z}^{186}$
$15:\mathbb{Z}^{589}$ $15:\mathbb{Z}^{233}$ $14:\mathbb{Z}^{20}$ $13:\mathbb{Z}^{604}$
$15:\mathbb{Z}^{40}$
Table 3. Homology of $4$-cut complexes $\Delta_{4}(C_{n}^{p})$
We have proved that for $n\geq 9=4\cdot 2+1$, the $3$-cut complex $\Delta_{3}(C_{n}^{2})$ is shellable and is homotopy equivalent to ${\binom{n-4}{2}-9}=\frac{n^{2}-8n-n+2}{2}$ wedge of spheres of dimension equal to the dimension of $\Delta_{3}(C_{n}^{2})$. It is easy to check that for $p\geq 3$, if $n\leq 2p+2$, then $\Delta_{3}(C_{n}^{p})=\emptyset$. Based on our calculation given in Table 2, we make the following conjecture for $n\geq 2p+3$.
Conjecture 4.2.
Let $p\geq 3$. Then $\Delta_{3}(C_{n}^{p})$ satisfies the following.
  1. (i)
    For $2p+3\leq n\leq 4p$, $\Delta_{3}(C_{n}^{p})$ is not shellable. Moreover,
    $\Delta_{3}(C_{n}^{p})\simeq\begin{cases}\mathbb{S}^{p-1},&\text{ if $n=2p+3$ and $p$ is even},\\ \mathbb{S}^{p},&\text{ if $n=2p+3$ and $p$ is odd,}\\ \bigvee\limits_{\alpha}\mathbb{S}^{n-8},&\text{ if $2p+4\leq n\leq 3p-1$,}\\ \bigvee\limits_{\beta}\mathbb{S}^{n-6},&\text{ if $3p\leq n\leq 4p-1$,}\\ \bigvee\limits_{p-1}\mathbb{S}^{n-5},&\text{ if $n=4p$,}\end{cases}$
    where $\alpha$ and $\beta$ depend on $n$ and $p$.
  2. (ii)
    For $n\geq 4p+1$, $\Delta_{3}(C_{n}^{p})$ is shellable and
    $\Delta_{3}(C_{n}^{p})\simeq\bigvee\limits_{\frac{n^{2}-4np-n+2}{2}}\mathbb{S}^{n-4}.$
$p$ $n$ 9 10 11 12 13 14 15 16 17 18 19
2 $1:\mathbb{Z}$ $3:\mathbb{Z}$ $5:\mathbb{Z}^{56}$ $6:\mathbb{Z}^{152}$ $7:\mathbb{Z}^{300}$ $8:\mathbb{Z}^{505}$ $9:\mathbb{Z}^{776}$ $10:\mathbb{Z}^{1125}$ $11:\mathbb{Z}^{1565}$ $12:\mathbb{Z}^{2110}$ $13:\mathbb{Z}^{2775}$
$4:\mathbb{Z}^{10}$
3 $3:\mathbb{Z}$ $4:\mathbb{Z}^{39}$ $5:\mathbb{Z}^{15}$ $7:\mathbb{Z}^{41}$ $8:\mathbb{Z}^{62}$ $9:\mathbb{Z}^{64}$ $10:\mathbb{Z}^{68}$ $11:\mathbb{Z}^{72}$ $12:\mathbb{Z}^{76}$
$6:\mathbb{Z}$ $8:\mathbb{Z}^{56}$ $9:\mathbb{Z}^{165}$ $10:\mathbb{Z}^{381}$ $11:\mathbb{Z}^{732}$ $12:\mathbb{Z}^{1213}$ $13:\mathbb{Z}^{1825}$
$7:\mathbb{Z}^{13}$
4 $3:\mathbb{Z}^{2}$ $5:\mathbb{Z}^{29}$ $7:\mathbb{Z}^{228}$ $8:\mathbb{Z}^{331}$ $9:\mathbb{Z}^{155}$ $10:\mathbb{Z}^{23}$ $11:\mathbb{Z}^{19}$
$4:\mathbb{Z}$ $10:\mathbb{Z}$ $11:\mathbb{Z}^{163}$ $12:\mathbb{Z}^{360}$
$11:\mathbb{Z}^{68}$ $12:\mathbb{Z}^{180}$ $13:\mathbb{Z}^{399}$
5 $5:\mathbb{Z}$ $6:\mathbb{Z}^{68}$ $8:\mathbb{Z}^{271}$ $9:\mathbb{Z}^{56}$ $11:\mathbb{Z}^{1426}$
$7:\mathbb{Z}$ $10:\mathbb{Z}^{606}$ $13:\mathbb{Z}^{19}$
6 $5:\mathbb{Z}^{2}$ $7:\mathbb{Z}^{8}$ $9:\mathbb{Z}^{705}$
$:\mathbb{Z}$ $8:\mathbb{Z}^{78}$ $10:\mathbb{Z}$
Table 4. Homology of $5$-cut complexes $\Delta_{5}(C_{n}^{p})$
$p$ $n$ 10 11 12 13 14 15 16 17 18 19
2 $1:\mathbb{Z}$ $3:\mathbb{Z}$ $5:\mathbb{Z}^{136}$ $6:\mathbb{Z}^{402}$ $7:\mathbb{Z}^{855}$ $8:\mathbb{Z}^{1537}$ $9:\mathbb{Z}^{2507}$ $10:\mathbb{Z}^{3841}$ $11:\mathbb{Z}^{5630}$ $12:\mathbb{Z}^{7979}$
$4:\mathbb{Z}$
3 $2:\mathbb{Z}$ $4:\mathbb{Z}^{64}$ $5:\mathbb{Z}^{44}$ $6:\mathbb{Z}^{3}$ $8:\mathbb{Z}^{177}$ $9:\mathbb{Z}^{205}$ $10:\mathbb{Z}^{216}$ $11:\mathbb{Z}^{228}$
$3:\mathbb{Z}^{2}$ $6:\mathbb{Z}$ $7:\mathbb{Z}^{91}$ $9:\mathbb{Z}^{500}$ $10:\mathbb{Z}^{1258}$ $11:\mathbb{Z}^{2603}$ $12:\mathbb{Z}^{4635}$
$7:\mathbb{Z}^{28}$ $8:\mathbb{Z}^{150}$
4 $3:\mathbb{Z}^{2}$ $5:\mathbb{Z}^{58}$ $6:\mathbb{Z}^{4}$ $8:\mathbb{Z}^{917}$ $9:\mathbb{Z}^{772}$ $10:\mathbb{Z}^{246}$
$4:\mathbb{Z}$ $6:\mathbb{Z}^{5}$ $7:\mathbb{Z}^{409}$ $10:\mathbb{Z}^{34}$ $11:\mathbb{Z}^{180}$ $11:\mathbb{Z}^{342}$
$12:\mathbb{Z}^{570}$
5 $4:\mathbb{Z}^{4}$ $6:\mathbb{Z}^{101}$ $7:\mathbb{Z}^{2}$ $9:\mathbb{Z}^{476}$
$5:\mathbb{Z}$ $8:\mathbb{Z}^{451}$ $10:\mathbb{Z}^{1159}$
$9:\mathbb{Z}^{9}$
6 $5:\mathbb{Z}^{2}$ $7:\mathbb{Z}^{135}$
$6:\mathbb{Z}$ $8:\mathbb{Z}^{20}$
Table 5. Homology of $6$-cut complexes $\Delta_{6}(C_{n}^{p})$
Examining Tables 3-5 for $k=4,5,6$, we observe that for $p\geq 3$, the complexes $\Delta_{k}(C_{n}^{p})$ have non trivial homology in dimensions lower than the dimension of $\Delta_{k}(C_{n}^{p})$. These findings raise the question of whether the complexes $\Delta_{k}(C_{n}^{p})$, if nonvoid, are not shellable for $k\geq 4$ and $p\geq 3$.
1.3 focuses on the shellability of the $k$-cut complex $\Delta_{k}(C_{n}^{2})$ of the squared cycle graphs for $k\geq 3$. In Theorem 1.4, we have proved that for $n\geq 9=2\cdot 3+3$, $\Delta_{3}(C_{n}^{2})$ is homotopy equivalent to wedge of ${\binom{n-4}{2}-9}=\binom{n-1}{3-1}-(2^{3-1}-1)n$ spheres of dimension $n-4$. Using the data in Tables 3-5, we extend this discussion with the following conjecture about the homotopy type of $\Delta_{k}(C_{n}^{2})$ for $k\geq 4$.
Conjecture 4.3.
For $k\geq 4$ and $n\geq 2k+3$,
$\Delta_{k}(C_{n}^{2})\simeq\bigvee\limits_{{n-1\choose k-1}-(2^{k-1}-1)n}\mathbb{S}^{n-k-1}.$